OpenStudy (anonymous):

which is rational and which is irational squareroot 10, 1 and a half, negative0.7, circumference, squareroot 2, 5 over 3, and negative 1 fourth

OpenStudy (anonymous):

also can you round the number to a decimal

OpenStudy (anonymous):

it not circumference it pi

Directrix (directrix):

A rational number is a number that can be written as the ratio of two integers. square root 10 --- Irrational 1 and a half = 1.5 = 3/2 ---- Rational negative0.7 = -.7 = (-7)/10 --- Rational (circumference)--> pi ---- Irrational square root 2 ---- Irrational 5 over 3 = 5/3 --- Rational negative 1 fourth = -1/4 --- Rational

Directrix (directrix):

>>can you round the number to a decimal<< Which number or numbers do you want to round to a decimal?

OpenStudy (anonymous):

square root 10, 1 and a half ,negative0.7 ,pi , square root 2 , 1 5 over 3 , negative 1 fourth